40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 26 21:01:26 2019, gautam, Update, LSC, POX and POY locking 
    Reply  Thu Jun 27 14:28:12 2019, gautam, Update, LSC, POX and POY locking 
       Reply  Tue Jul 2 17:34:54 2019, gautam, Update, LSC, Irides opened up on EY table 
          Reply  Tue Jul 9 18:06:05 2019, gautam, Update, LSC, Y-arm ASS in a workable state ditherFreqs.pdftransRenorm.png
Message ID: 14720     Entry time: Tue Jul 2 17:34:54 2019     In reply to: 14705     Reply to this: 14738
Author: gautam 
Type: Update 
Category: LSC 
Subject: Irides opened up on EY table 

In preparation for the ASS debugging, I decided to check out the beam path on the EY table. In order to be able to do this, I had to setup the POY locking to trigger on AS110 instead of TRY (as is usual for this kind of debugging). Then I could poke an IR card in the beam path without destroying the lock.

There are two irides in the beam path immediately between the vacuum window and the harmonic separator that splits off the IR and green beams. I found that the beam was in fact getting clipped on both of them. It was also somewhat off center on a 2" beamsplitter that sends half of the light to the QPD (currently decommissioned). The purpose of these irides are (I think) to eliminate some ghost reflections of the green beam and also the Oplev beam. I opened up the irides until I felt that there wasn't any more clipping of the IR beam, but the appropriate ghost beams were still getting caught.

I also re-aligned the beam onto the TRY Thorlabs PD so as to better center it on the active area. In summary, the result of this work was that the TRY level went from ~0.6 to ~0.93. There may still be some scope for optimizing this - I tried running the Y-arm ASS scripts, and already, the loops don't run away any more. I'll do the systematic analysis of the servo anyways. But given that the IMC Trans lev el used to be ~15,500 counts and is now ~14,500 counts, I think ~7% drop in TRY level is in line with what we "expect" (assuming the pre-power-degradation TRY level was 1.000).

Note that these irides were installed (I think) by Yuki, and so cannot explain the ASS anomalies of July 2018 (i.e. it does not exonerate in-vacuum clipping of the beam, as Koji had already verified that the in-air path was clean back then).

ELOG V3.1.3-